OpenStudy (anonymous):

The base of a solid is the circle x2 + y2 = 4. Find the volume of the solid given that the cross sections perpendicular to the x-axis are squares.

OpenStudy (campbell_st):

well the problem is \[V = \pi \int\limits_{0}^{2} y^2 dx\] or \[V = \pi \int\limits_{0}^{2} 4 - x^2 dx\]

OpenStudy (anonymous):

The answer choices are 131/3 , 128/3, 137/3, 134/3, 140/3 and teh answer i get when i do taht integration is 16pi/3

OpenStudy (anonymous):

and 16pi/3 = 16. 75 and none of the asnwer choices equal that